OpenStudy (anonymous):

(6, -4): 2x-7y=9. just need the answer please.

OpenStudy (anonymous):

what u want? specify...

OpenStudy (anonymous):

it says write an equation of the line containing the given point and parallel to the given line. y=

OpenStudy (anonymous):

(y+4) =2/7(x-6) => y = 2/7x - 12/7 -4 => y = 2/7 x - 40/7

OpenStudy (anonymous):

Thank you all that got me thru it.